### v1

# Overestimation of melting temperatures calculated by first-principles molecular dynamics simulations

Abstract
Although the melting temperature, $T_{m}$, of a solid can be calculated based on first-principles molecular dynamics (FP-MD) simulations, systematic assessments of the accuracy of the resulting values have not yet been reported. FP-MD simulations require significant computational resources and hence an examination of the effect of cell size on convergence is difficult. In addition, calculation of the energy of a liquid is not a trivial problem because of energy dissipation effects. The present work attempts to resolve these problems, and thus allow the accuracy of $T_{m}$ values obtained from FP-MD simulations to be assessed for typical semiconductors, metals, and oxides. With the exception of Si, the $T_{m}$ value was overestimated in all cases. This overestimation can be reduced by increasing the cell size, although the convergence is slow unless the potential is very shallow. For oxides, this overestimation may not be removed by increasing the cell size. The LDA/GGA error of overbinding affects the melting enthalpy and thereby $T_{m}$. In order to fully capture the energy dissipation nature of liquids, adiabatic MD simulations are required, and such simulations have been performed in the present study.
I Introduction
Numerous successes of density-functional theory (DFT) calculations in recent materials research promote the status of DFT calculations to almost an exact science. On this basis, first-principles (FP) molecular dynamics (MD) simulations, which are based on DFT, would be expected to accurately predict the melting temperature ($T_{m}$). Although there are many studies to apply FP-MD simulations to liquids Stich et al. (1991); Kresse and Hafner (1994); Chelikowsky et al. (2001); Massobrio et al. (2015), little studies focus $T_{m}$ as the main topics. The primal interests of these studies are the structural properties of liquids, such as the radial distribution function (RDF), the transport properties, and reaction kinetics. Accordingly, little attention was paid for the accuracy of $T_{m}$ values obtained from FP-MD simulations. Although FP calculations do not involve fitting parameters, control parameters exist in real FP codes and the results are to some extent affected by these parameters. Thus, it is necessary to examine the effects of such parameters on convergence, even though this often requires high computational loads
Many researchers studying melting phenomena continue to use classical MD simulations with model potentials, because of needs for prolong and large-scale simulations. The success of such simulations largely depends on the accuracy of the model potentials, which are usually tested by comparing selected properties of the material between calculation and experiment. As such, structural parameters are most commonly used. Phonon calculations using the force constant model can be looked upon as an example of construction of model potential. Often only few parameters can give a good agreement with an experimental phonon spectrum Bruesch (1982). However, agreement in one case does not guarantee agreement with respect to other properties. As an example, Porter et al. demonstrated that different models that produce almost identical phonon spectra predicted completely different thermal expansion values Porter et al. (1997). This inconsistency could be improved by increasing the number of parameters in a model. Even so, in the case of phenomena for which the harmonic approximation provides an inaccurate prediction, it remain unclear whether a suitable model potential can be constructed based on a reasonable number of parameters. Melting involves increased uncertainty, in that no reliable model superior to the existing Lindemann rule is available Grimvall and Sjödin (1974); Granato et al. (2010).
Reviewing previous studies concerning the melting behavior of silica-based materials illustrates the challenge associated with devising satisfactory models of potential for silica. In fact, there have been many attempts to devise potentials for silica so as to accurately reproduce their physical properties. By constructing interatomic potentials of Si-O pairs in order to match the experimental values, Yamahara et al. were able to correctly reproduce the negative thermal expansion of $\beta$-cristobalite and the $\alpha$-$\beta$ transition of cristobalite of approximately 500 K Yamahara et al. (2001). However, the same potentials produced a $T_{m}$ value for $\beta$-cristobalite about 5000 K, which is much higher than the experimental value of 1996 K. The same MD simulations also largely overestimated for the glass transition temperature, $T_{g}$, giving a value of 3800 K compared with the experimental result of 1480 K. Predicting the glass transition temperature of silica is an especially challenging task. Although many MD simulations have been reported Vollmayr et al. (1996); Yamahara et al. (2001); Kuzuu et al. (2004); Takada et al. (2004); Carré et al. (2008); Geske et al. (2016); Niu et al. (2018), the calculated values of $T_{g}$ are scattered over a wide range of temperatures from 2000 to 4000 K. Takada, et al. report that popular potentials TTM and BKS reproduce the structures of silica well, but provide unsatisfactory $T_{g}$ Takada et al. (2004). Here it is helpful to note that the structural parameters that are matched with experiment when potentials are constructed are the time-averaged structures, such as the bond length. RDF is in this category. The distinction between the solid and liquid states is smeared out in those averaged quantities. The time correlation becomes more important for liquids, as discussed in this paper. We recall the often-claimed leaning that real potentials have highly non-analytic forms, which explains why the DFT approach is so valuable.
DFT should provide, in principle, the correct answer to this problem, because the potentials are constructed from very fundamental principles of physics. However, in practice, use of DFT does not automatically guarantee the correctness of the results. First, even if the potential were exact, the final result will only be correct in the case that the MD simulations are properly set up. The most serious problem in this respect is imposing the periodic-boundary conditions. Melting involves a break-up of the periodicity of crystal, and so melting simulations are performed using the largest supercells that the computational resource permits. In reality, the practical limit is a few times the size of the primitive unit cell. The imposing the periodic-boundary conditions causes several spurious effects. The effects of finite-size supercell are well studied for the defect physics Castleton et al. (2009); Freysoldt et al. (2014). However, different types of consideration are required for the melting phenomenon, which has not been studied in depth to date.
Second, the currently available DFT potentials are still approximations. The exact energy functional in DFT is not known, and will likely not be found in the near future. The overbinding problem of LDA/GGA is a well-known problem with these approximations Jones and Gunnarsson (1989); Parr and Yang (1989). However, the influence of the overbinding on $T_{m}$ is not well studied not (a). This requires calculation of the internal energy of liquid. Although calculation of energy itself is easy in DFT, it is not a trivial task for liquids. Liquids exist only at finite temperatures, and hence the calculation of $E$ without specifying $T$ has no sense. We need to determine both $E$ and $T$ as a set. Thus the problem is equivalent to calculating specific heat $C$, for which the standard theory is lacked for liquids Granato (2002); Trachenko and Brazhkin (2011); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Proctor (2020); Baggioli and Zaccone (2021). Determination of the relationship between $E$ and $T$ is critical. Energy dissipation (atom relaxation) processes play an important role in determining the $E-T$ relation, and careful consideration of this effect is required for the set up of MD simulation. This is the topic that the present authors have recently established for the glass transition Shirai et al. (2022, 2023). By utilizing the method of calculating $E$ and $T$, we are able to discuss the effect of overbinding on $T_{m}$.
In this paper, the accuracy of $T_{m}$ values obtained by FP-MD simulations is assessed by the method employed in previous studies Shirai et al. (2022, 2023). In Sec. III.1, we report that the $T_{m}$ values obtained by FP-MD simulations are almost always overestimated by examining various materials. The reason for the overestimation is not unique, and depends on material. Further analysis is made for selected materials in the remaining part of Sec. III. These examples are chosen from semiconductors, metals, and oxides. This selection hopefully provides a useful guidance for further study. At present, computationally efficient methods for the determination of $T_{m}$ are available, including thermodynamic integration Frenkel and Smit (1996); Sugino and Car (1995); Cheng et al. (2019). However, the primitive method of adiabatic MD simulations remains important, because the accuracy of calculations is inseparable from an appropriate understanding of the liquid state thermodynamics, as explained in the next section.
II Methods
II.1 Simulation of melting
This work employed both dynamical and thermodynamical methods to determine $T_{m}$ on the basis of MD simulations. The most commonly used method may be the calculation of diffusion constant $D$. $T_{m}$ is determined as the temperature at which a finite value $D$ appears as a solid is heated. The calculation of $D$ is a quite standard method Frenkel and Smit (1996), and thus no explanation is needed here. The other method is the calculation of the internal energy $U$, because $U(T)$ exhibits an abrupt change at $T_{m}$. This, in turn, provides the latent heat associated with the melting enthalpy $H_{m}$. Because of the importance of energetics, we focus more on this $U$ method. Recently, the authors developed a method of calculating the specific heat associated with the glass transition Shirai et al. (2022, 2023). This method is advantageous in that it provides the specific heat and other thermodynamic functions for both solids and liquids with an equal level of accuracy. This method is described below in order to emphasize the full thermodynamic treatment of the liquid state.
II.1.1 Energy vs. temperature
Let us consider an MD simulation for a system composed of $N$ atoms. The $j$th atom having mass $M_{j}$ is at position ${\bf R}_{j}(t)$ with velocity ${\bf v}_{j}(t)$ at time $t$. On the Born-Oppenheimer approximation, the total energy, $E_{\rm tot}$, of a system is given by the sum of the kinetic energy, $E_{\rm K}$, and the potential energy, $E_{\rm P}$, as
$E_{\rm tot}(t)\equiv E_{\rm P}+E_{\rm K}=E_{\rm gs}(\{{\bf R}_{j}(t)\})+\frac{1}{2}\sum_{j}M_{j}v_{j}(t)^{2},$ (1)
where $E_{\rm gs}(\{{\bf R}_{j}(t)\})$ is the ground state energy of the electrons, including ion-ion potentials, for the instantaneous positions $\{{\bf R}_{j}(t)\}$. The internal energy in thermodynamics is defined at equilibrium and is given by the time average of the microscopic total energy $E_{\rm tot}(t)$ of the system as
$U=\overline{E_{\rm tot}(t)}=\overline{E_{\rm gs}(\{{\bf R}_{j}(t)\})}+\frac{1}{2}\sum_{j}M_{j}\overline{v_{j}(t)^{2}}.$ (2)
For solids, the constituent atoms fluctuate around their equilibrium positions. The instantaneous position of the $j$th atom can be expressed by the sum of the equilibrium position, $\bar{\bf R}_{j}$, and a small displacement from this position, $\bar{\bf u}_{j}$, as ${\bf R}_{j}(t)=\bar{\bf R}_{j}+{\bf u}_{j}(t)$. The ground-state energy, $E_{\rm gs}(\{{\bf R}_{j}(t)\})$, can be expanded in the Taylor series in terms of this displacement. The part of time dependent term together with $E_{\rm K}$ constitutes the phonon energy and its time average $E_{\rm ph}$ is given by
$E_{\rm ph}=\frac{1}{2}\sum_{i,j}\overline{{\bf u}_{i}(t)\cdot{\bf D}_{ij}\cdot{\bf u}_{j}(t)}+\frac{1}{2}\sum_{j}M_{j}\overline{v_{j}(t)^{2}},$ (3)
where ${\bf D}_{ij}$ is the dynamic matrix between the $i$th and $j$th atoms Born and Huang (1969). The first term of the right-hand side of Eq. (3) is referred to $E_{\rm P,vib}$ herein. The remaining part, $E_{\rm st}(\{\bar{\bf R}_{j}\})$, in the expansion $\overline{E_{\rm gs}(\{{\bf R}_{j}(t)\})}$ is constant with respect to time and is referred to as the structural energy. Thus, the internal energy $U$ consists of the following terms,
$U\equiv U(T,V,\{\bar{\bf R}_{j}\})=E_{\rm st}(\{\bar{\bf R}_{j}\})+E_{\rm ph}(T)+E_{\rm te}(V).$ (4)
In Eq. (4), three translation and three rotational displacements of the entire system have been removed from the term $E_{\rm st}(\{\bar{\bf R}_{j}\})$. The contribution of these displacements is extracted to form the last term $E_{\rm te}(V)$, which is referred to as the thermal expansion energy. Therefore, the structural energy $E_{\rm st}$ means the energy change by the internal coordinates with fixed lattice parameters. Figure 1 summarizes the relationships between $U$ and these components.
Figure 1: The relationships between various energies in adiabatic MD simulations. As $T$ increases, the bottom of the adiabatic potential is raised by $E_{\rm st}$.
By diagonalizing the dynamic matrix, $E_{\rm ph}$ can be calculated as
[…]
where $\hbar$ is the Planck constant, $\bar{n}(\omega)$ is the Bose occupation number, and $g(\omega)$ is the phonon DOS. Because atom motions are treated classically in conventional MD simulations, the kinetic energy of atoms in these simulations does not obey the Bose-Einstein statistics but instead obeys the classical equipartition law of energy,
$\overline{E_{\rm P,vib}}=\overline{E_{\rm K}}=\frac{3}{2}Nk_{\rm B}T,$ (6)
where $k_{\rm B}$ is Boltzmann’s constant. In order to fix this problem, first this classical term in $\overline{E_{\rm gs}(t)}$ is removed from the total energy, giving the structural energy $E_{\rm st}$,
$E_{\rm st}=\overline{E_{\rm gs}(t)}-\overline{E_{\rm K}}=\overline{E_{\rm gs}(t)}-\frac{1}{2}E_{\rm ph}.$ (7)
Then, the phonon energy $E_{\rm ph}$ is evaluated through Eq. (5). The phonon DOS, $g(\omega)$, is obtained from the power spectrum of atom velocities in an MD run Payne et al. (1992), whereas the temperature in the Bose occupation number is determined by the kinetic energy of atoms via Eq. (6). Finally, $E_{\rm tot}$ is recalculated by adding this phonon energy $E_{\rm ph}$ to $E_{\rm st}$, recovering the Bose-Einstein statistics. In practice, this modification of $E_{\rm tot}$ is important only at very low temperatures.
For liquids, the atoms do not have unique equilibrium positions and hence the expansion (3) does not make sense. Despite this, we can operationally define a “phonon” energy for a liquid in a manner similar to that described above in Eq. (5). This is possible because the frequency spectrum of atom velocities, $f_{v}(\omega)$, can be obtained from MD simulations. In fact, deriving the phonon dispersion in this manner is a practice in the inelastic neutron scattering analysis Egelstaff (1992); Copley and Rowe (1974); Smith et al. (2017). Observation of even a Brillouin-zone-like structure is surprising Suck (2005). The frequency spectrum of atom velocities $f_{v}(\omega)$, however, does not distinguish scattering from vibrational motions. The idea of $f_{v}(\omega)$ applies even to an ideal gas, in that collisions create oscillatory behavior. This phonon-like scenario is often used in modeling liquids, which often gives a good description of the specific heat $C_{p}$ Wallace (1997, 1998); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Baggioli and Zaccone (2021). We note, however, that the specific heats of many metal liquids are, from beginning, close to the classical limit of the equipartition law of energy, $C_{v}=3k_{\rm B}$ Wallace (2002), which is insensitive to the mode of atom motions. Throughout this paper, the specific heat per atom is understood by using the unit $k_{\rm B}$. Upon closer observation, it will be seen that the phonon-like model described above is only an approximation, as discussed in Sec. III.2. A good example is water, for which $C_{p}\cong 3k_{\rm B}$ even though H-related vibrations cannot be thermally activated at the $T$ range of relevance. Liquid Se has a large $C_{p}$ more than 4 $k_{\rm B}$ Shu et al. (1980). In many metallic liquids, $C_{v}$ near $T_{m}$ falls in a range 3.0 to 3.4 $k_{\rm B}$, which already exceeds the classical limit (Ref. Wallace (2002), p. 242). Some authors ascribe this excess specific heat to the anharmonic effect of phonons Wallace (1998); Bolmatov et al. (2012); Trachenko and Brazhkin (2016); Proctor (2020); Baggioli and Zaccone (2021). But, this excess specific heat increases as $T$ approaches $T_{m}$ from the high temperature side. The anharmonic effect of phonons is expected to be reduced upon lowering $T$ Leibfried and Ludwig (1961). Moreover, recent experiments on supercooled liquids show that the increasing behavior of the excess specific heat as $T$ decreases is continuously extended in the range $T<T_{m}$, namely, $dC_{p}/dT\approx$ const with a negative constant value: Perepezko and Paik (1984) for metallic liquids, Rhim and Ishikawa (2000); Li et al. (2004) for melton Ge. Therefore, it is inappropriate to ascribe the increasing behavior of the excess specific heat with decreasing $T$ to the anharmonic effect of phonons.
In spite of facile success of phonon theory for liquids, the theory eventually fails to accurately describe liquid states not (b). Around the transition temperature of phase transition, the contribution of $E_{\rm st}$ plays an important role in the $T$ dependence on $U$, as shown by previous MD simulations Shirai et al. (2022, 2023). Recent studies on highly viscous glass-forming materials, show that the configuration contribution—this corresponds to $E_{\rm st}$ in this paper—dominates the excess entropy in the supercooled liquid state Smith et al. (2017); Alvarez-Donado and Antonelli (2020); Han et al. (2020). The contribution from $E_{\rm st}$ must therefore be taken into account. In the present method, the total energy $E_{\rm tot}$ includes both the components $E_{\rm st}$ and $E_{\rm ph}$. In the sprit of DFT, the total energy is the most reliable quantity.
The difficulty in calculating the thermodynamic functions of liquids lies essentially in that there is no eigenstate for liquids. In standard applications of statistical mechanics, the presence of eigenstates $\{i\}$ is presumed, when the partition function, $Z=\sum_{i}e^{-\varepsilon_{i}/k_{\rm B}T}$, is evaluated. In the phonon model, phonon-like excitations is envisaged for liquids. But this phonon is soon destroyed after creation. The energy conservation requires that the destruction of a phonon is compensated for by creation of new phonons Leibfried and Ludwig (1961). The liquid state can therefore be considered as a steady state achieved by the dynamic balance between incessant creation and destruction of phonon-like excitations. In an adiabatic simulation the total energy of the system is conserved and the temperature is adjusted to achieve this dynamic balance. Such a system with constant $NVE$ is usually referred to as the microcanonical ensemble, but here the name adiabatic MD simulation is used. The importance of the energy dissipation in determining $T$ is demonstrated by the frequency dependence of the specific heat in the glass transition Birge and Nagel (1985); Nielsen and Dyre (1996); Hentschel et al. (2008). This cannot be explained by the phonon theory. By construction, this energy dissipation process is fully reflected in MD simulations because of the equations of atom motions, and hence the correct $U-T$ relationship can be automatically obtained by MD simulations. However, if the viscose terms coupled to the thermostat are introduced into the MD simulation, the the energy dissipation process is modified by this artificial coupling between the system and the thermostat, destroying the intrinsic relationship between $U$ and $T$. Adiabatic ($NVE$) simulations are therefore necessary to obtain the correct $U-T$ relationship .
II.1.2 Construction of MD simulations
In this study, the internal energy $U$ must be evaluated over a wide range of $T$ to cover from the solid to liquid states. To ensure that equilibrium was established at each $U$, an adiabatic MD run was performed for a fixed $E$ and repeated by changing the total energy. In adiabatic MD simulations, the control parameter is the initial temperature $T_{\rm in}$, at which the initial velocities of atoms are given through Eq. (6). In each single MD run, the simulation was continued until equilibrium was established. It is noted that there are two kinds of equilibrium conditions Shirai (2020). The one is the equilibrium condition with respect to temperature (thermal equilibrium). The equilibrium temperature is determined by
$\frac{3}{2}k_{\rm B}T=\frac{1}{2}\langle M_{j}\overline{v_{j}(t)^{2}}\rangle_{j}.$ (8)
Here, the brackets denote the particle average, whereas the bar indicates the time average. Thermal equilibrium is reached when the time average of $E_{\rm K}(t)$ converges. In the following, $T$ in all figures indicates this equilibrium temperature. The time required to reach this condition is referred to as the thermal relaxation time, $\tau_{T}$. For most of solids, $\tau_{T}$ is quite short, on the order of 0.1 ps. For liquids, $\tau_{T}$ becomes longer but is still less than 1 ps. Thus, in most cases the relaxation to equilibrium is controlled by the other condition, namely, the mechanical (structural) relaxation. The time required to this relaxation is referred to as the structural relaxation time, $\tau_{M}$. The time evolution of the particle-averaged displacements,
$\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}=\frac{1}{Nt_{\rm SM}}\sum_{j}^{N}\int\left(R_{j}(t-t_{0})-R_{j}(t_{0})\right)^{2}t_{0},$ (9)
is used to determine whether structural equilibrium is reached, where $t_{\rm SM}$ is the simulation time. For solids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ is constant with respect to $t$. In contrast, for liquids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ shows a linear dependence on $t$ over the entire simulation time $t_{\rm SM}$. For the latter case, the slope, $D^{\prime}=\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}/t$, gives the diffusion coefficient. Although the true diffusion coefficient, $D$, is given by $D=(1/6)\lim_{t\rightarrow\infty}\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}/t$, $D^{\prime}$ is used for the diffusion coefficient in the following. For the liquids treated in this study, $\tau_{M}$ was on the order of a few ps. Examples of the time development of $E_{\rm K}$ and $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ are provided in the Supplemental Material to Ref. Shirai et al. (2023).
II.2 Calculation conditions
While several codes were used for the present FP-MD simulations, the majority of calculations were performed by Phase/0 Yamasaki et al. (2019). Some were performed by a code written in-house (Osaka2k) or the VASP code VAS . All codes are pseudopotential methods using planewave expansion. The MD simulations employed time steps ranging from 0.72 to 1.2 fs and the total simulation time, $t_{\rm SM}$, varied from 2 to 10 ps, depending on the relaxation time. The volume of the crystal, $V$, was fixed at the experimental value, although normally the $V$ of the corresponding liquid phase is slightly larger than that of the crystal. The effect of volume change on $T_{m}$ is examined in subsection III.4.
III Results and discussion
III.1 Melting temperature
Melting simulations for various materials were performed by FP-MD simulation by scanning a wide range $T$ covering both the solid and liquid states. Table 1 summarizes the calculated $T_{m}$ together with experimental values for comparison. The potential type is indicated by NC (norm-conserving), US (ultrasoft), PAW (projector augmented waves) Martin (2004), along with the type of electron correlation energy functional (LDA or GGA). The cutoff energy $E_{\rm cut}$ for the planewave expansion is given in units of eV. In MD simulations, melting occurs over a finite range of $T$, as shown later. Here this $T$ range is referred to as the transition region. The calculated value of $T_{m}$ is then defined as the mid point of the transition region. The width, $\Delta T_{m}$, of the transition region is expected to approach zero as the size of supercell is increased. In the table, the deviation of the calculated value from the experimental value is indicated by the ratio, $r$, of the former to the latter. In all the cases studied, $r>1$ was obtained, meaning that the calculated values were always overestimated. It is our surprise to obtain large values of $r$ more than 2, despite the initial confidence in the DFT calcluations. Large overestimations of $r>2$ are found when the material contains oxygen atoms. Only for Si, the calculated value is in a good agreement with experiment.
In the table, the melting enthalpy $H_{m}$ is also listed in units of eV/atom. $H_{m}$ is obtained from the abrupt change in $E_{\rm st}$. In experiment, the change $H_{m}$ appears at a discrete temperature $T_{m}$, whereas, as noted, melting occurred over a finite width $\Delta T_{m}$ in the present MD simulations. Hence, the accuracy of $H_{m}$ depends on the sharpness of the transition region. Small cell sizes give less reliable $H_{m}$ values. The accuracy of $H_{m}$ highly depends on the accuracy of $T_{m}$. These two quantities are related each other by
$H_{m}=T_{m}\Delta S_{m},$ (10)
where $\Delta S_{m}$ is the entropy of melting. This relationship says that an error of 0.1 eV in $H_{m}$ causes an error of 1000K in $T_{m}$, provided $\Delta S_{m}/k_{\rm B}=1$. Because $\Delta S_{m}$ in many materials is of this order of magnitude, the accuracy of $H_{m}$ directly controls the accuracy of $T_{m}$. Again the calculation for Si shows a good agreement in $H_{m}$ with experiment. This confirms the internal consistency between calculated $T_{m}$ and $H_{m}$ for Si. The phonon energy, $E_{\rm ph}$, and thermal expansion energy, $E_{\rm te}$, can contribute $H_{m}$. However, the previous studies Shirai et al. (2022, 2023) showed that these two contributions can generally be ignored, unless there is a large difference in thermal expansion between the solid and liquid phases. Because the melting temperature of most materials is high, the specific heat of the phonon contribution near $T_{m}$ becomes already the classical limit of the Dulong-Petit law, $C_{v}=3k_{\rm B}$, and thus there is no reason that an abrupt jump in $E_{\rm ph}$ occurs at $T_{m}$.
Material $T_{m}$ (K) $H_{m}$ (eV/atom) Size Potentials
exp. calc. $r$ exp. calc. (atoms) type $E_{\rm cut}$ code
Si 1683 1750 1.04 0.52 0.41 512 NC+GGA 204 Phase
Na 371 460 1.24 0.026 0.015 128 NC+GGA 245 Phase
$\alpha$-quartz (SiO2) 1700 4500 2.64 0.033 0.13 72 US+GGA 408 Phase
Glycerol (C3H8O3) 291 635 2.18 0.013 0.010 56 US+GGA 408 Phase
B2O3 723 3000 4.15 0.051 0.18 60 PAW+GGA 340 Phase
H2O 273 720 2.63 0.021 0.03 96 NC+LDA 816 Osaka2k
As2Se3 645 1080 1.67 0.037 0.1 160 US+GGA 408 Phase
Table 1: Comparison of calculated $T_{m}$ to the experimental value for various materials. $r$ indicates the ratio of the calculated $T_{m}$ to the experimental value. The melting enthalpy $H_{m}$ is also listed. Size indicates the size of the supercell used in MD simulations. The last block indicates the used potentials. $E_{\rm cut}$ for planewave expansion is given in eV units. $\Gamma$-point $k$ sampling is used in all the listed data. Experimental data are from CRC (2011) with the exception of the data for quartz Richet et al. (1982).
Several causes for the error in the calculated $T_{m}$ are investigated. One important cause is the use of periodic boundary conditions. An effect of the periodic boundary conditions is elimination of the surface effects. Because melting first occurs at the surface of a material, it can be argued that elimination of the surface makes melting less likely, thus increasing $T_{m}$. This effect could contribute to the overestimation of $T_{m}$. However, based on the good agreement between the calculated and experimental $T_{m}$ values for Si, this cannot be the primary reason for the large overestimation of $T_{m}$ unless a very large surface energy is involved. The introduction of defects could have a similar effect on melting. However, in our experience for quartz Shirai et al. (2023), the introduction of a defect in a supercell did not affect $T_{m}$ within the numerical accuracy.
Another effect of the periodic boundary conditions is the creation of spurious reflections of atom movements, which yields an artificial barrier to such movements. For a given size of supercell, $L$, all long-wavelength phonons (those with wavelength $\lambda>L$) are eliminated. In a previous paper Shirai et al. (2023), by using a crude model, it is shown that this artificial barrier, $V_{0}$, is reduced only in a scale $1/L^{2}$, while the computational demands increase as $L^{3}$. The value of $V_{0}$ can be of the order of a few tens of eV with a cell size of $2\times 2\times 2$ for $\alpha$-quartz. However, to test this scaling, real calculations is needed by increasing the cell size. In practice, extending the size of supercells beyond those listed in Table 1 is not feasible. For a 512-atoms simulation, one month computation was required for a single temperature in our cluster machines (typically in an 8-core parallel configuration). Hence, the effect of supercell size can realistically only be examined by reducing the supercell size, even though simulations become unrealistic. In the following, the effect of supercell size is examined by taking one example from each of the categories of semiconductors, metals, and oxide compounds. All the detailed data listed in Table 1 are seen in the Supplemental Materials.
III.2 Silicon
The first example is Si. The cell sizes of $N=$8, 64, 216, and 512 atoms were examined. In Fig. 2, the structural energy, $E_{\rm st}$, and diffusion constant, $D^{\prime}$, are plotted as functions of $T$. Note that $E_{\rm st}$ is normalized by the number of atoms in the cell and is in units of K. This is convenient because the specific heat is immediately obtained in units of $k_{\rm B}$, making it easier to compare the obtained value to the classical limit $3k_{\rm B}$. This convention is used throughout this paper. In this figure, the data points are connected by lines to show the sequence of changing $T_{\rm in}$. As shown in Fig. 2, $E_{\rm st}$ is essentially constant with changes in $T$ up to $T=1000$ K, indicating that the crystal potential is well described by the harmonic approximation. Finite values of $D^{\prime}$ appear approximately $T=2000$ K, confirming the onset of melting. In accordance with the variation in $D^{\prime}$, $E_{\rm st}$ exhibits an abrupt increase.
Figure 2: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, of Si calculated by NC+GGA. The cell sizes ($k$ meshes) are $N=8$ ($4^{3}$), 64 ($2^{3}$), 216 (1), and 512 (1). $E_{\rm cut}=204$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin is set to the ground state energy of that cell.
Sigmoidal behaviors with negative slopes are observed in both $E_{\rm st}$ and $D^{\prime}$ around $T_{m}$. The negative slope in $E_{\rm st}$ against $T$ means a negative specific heat, which is thermodynamically unstable Lynden-Bell (1999); Landsberg and Pearić (1987); Michaelian and Santamaria-Holek (2007). The negative slope in $D^{\prime}$ against $T$ also means dynamical instability. This behavior was pointed out previously as the effect of finite-size cells Shirai et al. (2022, 2023). In a finite-size MD simulation with $N$ atoms, the kinetic energy has a fluctuation of $\Delta T/T=1/\sqrt{N}$. When the crystal melts, the energy distribution between the vibrational motions (in the solid portion) and translational motions (in the liquid portion) becomes unbalanced due to this energy fluctuation. At temperatures slightly below $T_{m}$, those atoms with sufficiently high kinetic energy (higher than $k_{\rm B}T_{m}$) begin to convert their vibrational motions to diffusing motions, decreasing the average kinetic energy $\langle E_{K}\rangle$ and increasing diffusion. In contrast, when this low-energy part of tge distribution are overpopulated, the system temperature is lowered. This makes these low-energy diffusing atoms being trapped in the crystal potential, leading to slight heating and a decrease in $D^{\prime}$. In this manner, an oscillatory behavior appears around $T_{m}$. In experiment, this sigmoidal behavior has been observed in clusters containing small numbers of atoms Schmidt et al. (2001). The transition region is narrowed as the size of the cell increases. On the limit $N\rightarrow\infty$, $\Delta T_{m}\rightarrow 0$ is reasonably expected. However, it is noticeable that $\Delta T_{m}$ of 300 K remains even for the cell size of 512 atoms. The convergence of the energy fluctuation is slow.
For Si, even though the convergence for the width $\Delta T_{m}$ is very slow, the calculated $T_{m}$, which is obtained by the mid point in the transition region, already agreed with the experimental value, when a cell size of $N=64$. Even a small cell of $N=8$ is not bad, aside from the $\Delta T_{m}$ value. From the viewpoint of construction of FP pseudopotential, Si is the easiest element Bachelet et al. (1982); Troullier and Martins (1991). The potential is not deep and there are no semi-core levels. This may be the reason why the good agreement is obtained for $T_{m}$.
For the liquid state, a decrease in $E_{\rm st}$ against $T$ can be seen in Fig. 2. This negative $T$ dependence is unrealistic. In contrast, the total energy, $E_{\rm tot}$, is an increasing function of $T$, as should be. A full description of the $T$ dependence of $E_{\rm tot}$ will be given elsewhere. Here, we only mention that this decreasing behavior of $E_{\rm st}$ is a consequence of applying Bose-Einstein statistics to the “phonons” of the liquid state. The decrease in $E_{\rm st}$ means that an overly large $E_{\rm ph}$ value was subtracted from $\overline{E_{\rm gs}(t)}$ in Eq. (7). This $T$ range ($T>T_{m}$) is already the classical regime, in which the Dulong-Petit law is expected to hold, if phonon theory is valid. All “phonon” modes are fully activated, which means that the thermal energy $(1/2)k_{\rm B}T$ is assigned to each mode $q$ of phonons in Eq. (5), for which the independent particle picture for phonons is assumed. In experiment, the $C_{v}$ of liquids decreases from approximately $3k_{\rm B}$ near $T_{m}$ to approximately $2k_{\rm B}$ near the boiling temperature $T_{b}$ (Ref. Wallace (2002), p. 244). From the phonon theory, it can be interpreted that the degree of phonon freedom, $f$, per atom is effectively reduced from $f=6$ at $T_{m}$ to $4$ at $T_{b}$. Therefore, applying Eq. (5) to the liquid state leads to an overestimation of the phonon energy, resulting in the decrease in $E_{\rm st}$. The independent particle picture is not valid for the “phonons” of liquids. This invalidity of independent particle picture for liquids can be understood by considering the extreme case of an idea gas, where $C_{v}$ has the limiting value $f=3$, which is usually interpreted as the freedom of purely translational motions. When the frequency spectrum, $f_{v}(\omega)$, of atom velocity is calculated for an idea gas, we have a “phonon” spectrum, $g(\omega)$. Even though there is no restoring forces, atom collisions create the oscillatory behavior, yielding $f_{v}(\omega)$. Obviously, in this case, the phonon energy, Eq. (5) should not contribute to the $C_{v}$ of the gas. On this basis, it is apparent that applying Bose-Einstein statistics to “phonons” obtained from the velocity distribution of particles is inappropriate for fluids. The decomposition in Eq. (4) is not valid for liquids. Only the total energy $E_{\rm tot}$ has a physical reality.
III.3 Na
The next example is metallic Na, having a BCC structure in the solid state.
Figure 3: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, of Na calculated by NC+GGA. Cell sizes are $N=16$ and 128. The $k$ meshes are $2^{3}$ for $N=16$, whereas are $1$ (red circles) and $2^{3}$ (blue circles) for $N=128$. $E_{\rm cut}=245$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy for that cell.
Cell sizes of 16 and 128 atoms were examined. In Fig. 3, the structural energy, $E_{\rm st}$, and diffusion constant, $D^{\prime}$, are plotted as functions of $T$. Let us first check the cell size dependence on $T_{m}$. When different sizes of supercells are compared, the equivalent $k$ meshes should be use. Thus, the data obtained by a $2^{3}$ $k$-mesh of 16-atom cell (crosses in the figure) should be compared with the data obtained by a $1$ $k$-mesh of 128-atom (red circles). Clearly, the $T_{m}$ value is lowered from 600 to 460 K by increasing the cell size from 16 to 128. Accordingly, we can expect further decrease in the $T_{m}$ value, approaching the experimental value 371 K, by increasing the cell size.
For metallic systems, the $k$ mesh can affect the total energy calculation largely. Hence, the dependence of the $k$ mesh was also examined for the 128-atom cell. However, there was only a marginal change in $T_{m}$ between $1$ (red circles) and $2^{3}$ (blue circles) meshes, as shown in this figure.
Although a non-negligible overlap between the core and valence electrons can require a correction in some cases Louie et al. (1982), there is no difficulty in constructing the pseudopotential for this case. Therefore, the primary issue causing the incorrect $T_{m}$ for Na appears to be the slow convergence against the cell size.
III.4 Silica
The last example is an oxide material, namely, crystalline $\alpha$-quartz. The authors previously used this crystal as the parent material for the formation of silica glass and found the overestimation of $T_{m}$ Shirai et al. (2023). The primitive unit cell of $\alpha$-quartz has the rhombohedral Bravais lattice comprising three SiO2 units. The data for $\alpha$-quartz listed in Table 1 represent the results of MD simulations performed by the Phase code. These calculations employed the US+GGA with 72-atom cell having dimensions of $2\times 2\times 2$. Unfortunately, both codes Phase and VASP failed to keep long time MD simulations with a smaller cell having dimension of $1\times 1\times 1$. Under these conditions, a drastic error propagations in atom forces occurred at some steps, leading to a break in the simulation. Osaka2k code using NC-LDA was found to be stable against the long run of MD simulations using this cell size.not (c) Hence, the cell-size dependence for $\alpha$-quartz was examined by osaka2k code with the NC-LDA potential, using $E_{\rm cut}$=544 eV for the planewave expansion. The results are shown in Fig. 4. Although this cutoff energy is insufficient to achieve convergence of the total energy for oxide materials, the value represents a compromise required to allow MD simulations with the NC potentials. As a consequence of this low cutoff energy, the potential of oxygen becomes effectively softer, leading to lowering $T_{m}$. This effect led to the reduced value of 3500 K in Fig. 4 compared with that of 4500 K in Table 1. Even so, the value of $T_{m}=3500$ K is still much higher than the experimental result of 1700 K.
Figure 4: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, for $\alpha$-quartz calculated by NC+LDA. $E_{\rm cut}=544$ eV. p1 indicates results obtained with a $1\times 1\times 1$ cell ($4^{3}$ $k$-mesh), while p8 indicates a $2\times 2\times 2$ cell ($1$). p8’ indicates the same cell with a volume expansion of 16.9 %. The experimental value of $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy of that cell. For p8’, it was set to that for the p8 cell.
As shown in Fig. 4, the onset of melting almost does not change upon increasing the cell size from 9 to 72 atoms. The calculated $T_{m}$ seems rather to be saturated at the large overestimated value. Similar results were obtained for $\beta$-cristobalite in that increasing the unit cell dimension from $1\times 1\times 1$ to $2\times 2\times 2$ did not alter the overestimated $T_{m}$ (see Supplemental Materials). Furthermore, the effect of volume expansion was also examined. The actual volume of the liquid silica is larger than that of the crystalline $\alpha$-quartz by 16.9 %. The same $2\times 2\times 2$ cell was assessed in conjunction with this volume expansion, referred to as the p8’ data in Fig. 4. Even with this large volume expansion, there was no discernible change in $T_{m}$. Although the possibility of reducing $T_{m}$ by further increasing the cell size cannot be excluded, it is unlikely that the large overestimation ratio, $r$, would be fixed by this change.
We rather consider that the overestimation of $T_{m}$ for $\alpha$-quartz is primarily due to the error of the potential. The systematic error of LDA is well known in that LDA overestimates the binding energy and accordingly underestimates the lattice parameters about a few % Jones and Gunnarsson (1989). The LDA error is a result of adapting the energy functional obtained for homogeneous electron systems to localized systems. Hence, the error generally becomes larger for strongly localized states. The underestimation of lattice parameters in LDA is fixed by GGA Jones and Gunnarsson (1989); Perdew et al. (1996); Zhang and Yang (1998); Perdew and Kurth (2003). However, the overestimate of the binding energy still remains in GGA, although the error is largely improved. Oxygen atom has the largest error. Patton et al. report that the overbinding $\Delta$ by GGA is 0.53 eV for O2 molecule Patton et al. (1997). For H2O, $\Delta$ is only 0.03 eV, but this amounts to a change in $T_{m}$ by 300 K. When calculating the formation energies of metal oxides, the formation energy for the oxygen part is typically adjusted by approximately 0.7 eV/(O atom) Wang et al. (2006); Jain et al. (2011); Mutter et al. (2020).
The cohesive energy, $E_{\rm coh}$, is the energy difference between an isolated atom and the solid state, representing the extrema of localized and extended states, respectively. Therefore, the LDA/GGA error appears in $E_{\rm coh}$ most significantly. Because $H_{m}$ is the energy difference between the solid and liquid states, a similar effect is expected. Here is, however, a subtle problem of the electron correlation in liquids. Normally, the electron system having a high atom density represents the extended state (solid state), while an electron system having a low atom density represents the localized state (atom state). However, the average density of a liquid is close to that of the solid for a given material: the difference is only a few %. There are even cases, such as Si, in which the liquid state is denser than the solid state. Therefore, the average density and average atom distance are not suitable to estimate the degree of localization of electrons for the case of liquid. Rather, in Secs. II.1 and III.2, we have seen that energy dissipation processes play important roles in describing the thermodynamic properties of liquids.
The energy dissipation can be described by the time correlation in a given pair of atoms (Egelstaff (1992), Chap. 9),
$G_{ij}({\bf r},t)=\overline{\delta({\bf r}+{\bf R}_{i}(t^{\prime})-{\bf R}_{j}(t^{\prime}+t))}.$ (11)
The bar indicates the time average with respect to $t^{\prime}$. The distance between atoms in a particular atom pair, $R_{ij}$, immediately increases with time after melting. This correlation function may be approximately presented by $G_{ij}({\bf r},t)\approx e^{-t/\tau_{ij}}\delta({\bf r}-\bar{d})$, where $\bar{d}$ is the average atom distance. The first factor $e^{-t/\tau_{ij}}$ is important. The relaxation time $\tau_{ij}$ is determined by the energy barrier between atoms $i$ and $j$. This corresponds to the saddle point of the energy landscape. In contrast, RDF is obtained differently. The correlation is evaluated first by taking the particle average and then by taking the time average, as
$\bar{G}({\bf r},t)=\overline{\langle\delta({\bf r}+{\bf R}_{i}(t^{\prime})-{\bf R}_{j}(t^{\prime}+t))\rangle_{i,j}}.$ (12)
This correlation, and accordingly, RDF reflect only the average atom distance, which corresponds to the minima in the energy landscape. The information of energy dissipation is completely missed, and thus the diffusing nature of liquid is masked in RDF. Agreement not only in the energy minima but also in the saddle points is needed to reproduce the thermodynamic properties of liquids. Consequently, even though the averaged density does not change significantly on melting, it is highly likely that a significant change occurs in the electron correlation at $T_{m}$. Thus, the LDA/GGA error also affects the energy barrier between the solid and liquid states, leading to an overestimation of $T_{m}$.
Lastly, it is worthy to mention a distinct feature of the liquid phase of silica. Normally, significant changes in $E_{\rm st}$ occur only at phase transitions: in this study, near $T_{m}$. However, as shown in Fig. 4, for the silica liquid, an apparent increase in $E_{\rm st}$ with increasing $T$ persists over the entire $T$ range examined. This implies that the silica liquid is comprised of numerous distinct metastable states, which are separated each other by energy barriers. This could explain why the silica liquid has the highest viscosity among various liquids. However, many things about the thermodynamic properties of silica liquid still remain unknown.
IV Conclusion
This work demonstrated systematic overestimations of the melting temperature in FP-MD simulations. Two important factors, the size dependence of the cell and the LDA/GGA error, were studied. The importance of these factors vary depending on the material under consideration. The convergence with respect to the cell size is not fast, unless the potential is very shallow. The significant finding of this study is demonstration of the effect of the LDA/GGA error on $H_{m}$ and $T_{m}$. This error cannot be removed by increasing the cell size.
Throughout this study, correct understanding thermodynamic nature of liquids is emphasized. There is no eigenstate for liquids, and instead the thermodynamic equilibrium of a liquid is established by the balance between incessant excitation and deexcitation in atom motions. The phonon model often hides this important nature of liquids. The total energy in adiabatic MD simulations describes all the process of energy relaxation, and hence gives the correct relationship between $U$ and $T$.
Acknowledgment
The authors thank Prof. Trachenko for discussing the specific heat of glasses and liquids. The authors thank FORTE Science Communications (https://www.forte-science.co.jp/) for English language editing.